# Scalable Data-Driven Basis Selection for Linear Machine Learning Interatomic Potentials

### v1

@@ IV Conclusion
V Acknowledgements
The work was supported by the New Frontiers in Research Fund through an NFRF Exploration Grant, and by the Natural Sciences and Engineering Research Council of Canada through an NSERC Discovery grant and an NSERC-NSF Collaboration on quantum science and artificial intelligence grant. CO is a partner in Symmetric Group LLP which licenses force fields commercially. The authors acknowledge helpful discussions with Gábor Csányi and Cas van der Oord.
-A Radial Basis Functions
@@ V Acknowledgements
$\Phi(\mathbf{y})=\frac{1}{2}\lambda||\mathbf{y}||_{2}^{2}-\mathbf{b}^{T}\mathbf{y},\quad\mathbf{g}(\mathbf{y})=\lambda\mathbf{y}-\mathbf{b}$
@@
The BPDual algorithm is an active-set method. The active set method focuses on identifying constraints that are active (i.e., binding) at each iteration. For an estimated solution y, a constraint is active if $\textbf{a}_{j}^{\top}\textbf{y}$ lies at one of its bounds (i.e. $\textbf{a}_{j}^{\top}\textbf{y}=\mathbf{1}$ or $-\mathbf{1}$). The active set $S$ is a set of indices representing the active constraints at each iteration, which forms a submatrix of A that has full column rank. For instance, for some permuation $\mathrm{P}$
@@
The estimate y is stationary if it satisfies the constraints $-\mathbf{1}\leq\textbf{A}^{\top}\textbf{y}\leq\mathbf{1}$ and the objective gradient $\mathbf{g}(\mathbf{y})$ is a linear combination of active constraint gradients. Specifically, there exists $x_{S}$ such that:
@@
[…]
@@
It’s not worthy that for efficient computations, only the upper triangular part of a QR factorization of $S$ is updated at each iteration using the QRupdate.jl [28] package.
@@
When $\alpha=1$, the solution $+\Delta\textbf{y}$ y minimizes $\phi(\textbf{y})$ on the current active set $S$. If improvement is possible by removing a constraint from $S$, the algorithm considers directions $\Delta\textbf{y}$ that move y away from an active constraint, maintaining feasibility. Consider an inequality constraint $q=S_{p}$ that is currently active at its lower bound, i.e., $a_{q}^{T}\textbf{y}=-\mathbf{1}$. To move away from this bound, we select a direction $\Delta\textbf{y}$ satisfying $S^{T}\Delta\textbf{y}=e_{p}$, ensuring $\Delta\textbf{y}$ is a feasible direction that moves away from the active lower bound. Observe that:
@@
[…]
-C Lasso Path
@@
As $\lambda$ decreases to $\bar{\lambda}$, it is also necessary to assess the impact on $\mathbf{y}$, particularly with respect to the constraints $-\mathbf{1}\leq A^{T}\mathbf{y}\leq\mathbf{1}$. Subtracting $\bar{\mathbf{y}}$ from both sides and using the definition of $\Delta\mathbf{x}$, we obtain:
@@
-D Orthogonal Matching Pursuit
[…]
-E BRR Posterior Derivation
@@
$\bm{\Sigma}=(\alpha\mathbf{I}+\lambda A^{\top}A)^{-1},$
@@
$\ln p(b\mid A,\alpha,\lambda)=$ $-\frac{m}{2}\ln(2\pi)+\frac{1}{2}\ln|\bm{\Sigma}|$
@@
This marginal log-likelihood is maximized with respect to $\alpha$ (or $\bm{\alpha}$ in the ARD case) and $\lambda$, typically using optimization methods such as BFGS or L-BFGS.
-F Truncated Singular Value Decomposition (TSVD) Post-Processing

### v2

@@ 4 Conclusion
For the datasets used in this study, OMP and ASP demonstrate similar accuracy-cost trade-offs on average. OMP, being a greedy algorithm that only adds elements to the active set, is somewhat more efficient in terms of wall-clock time. ASP, on the other hand, offers a more principled path-tracing approach, as it follows the exact $\ell_{1}$-regularized solution path rather than relying on greedy updates. Both outperform the other sparse optimization methods we benchmarked, achieving better results in terms of both cost and accuracy. Looking forward, OMP has the potential advantage that it can be more easily extended to nonlinear regression settings.
It’s noteworthy that our approach is not restricted to single-component systems. Because it builds on sparse regression applied to a chosen feature set, the framework can be readily transferred to other materials once a suitable basis is defined. This makes the method broadly applicable, from simple elemental systems to complex alloys and compounds. The key limitation is that for systems with many chemical species (say, more than 4) the added combinatorial complexity makes linear model inherently expensive. In this case one should apply our approach to a linear model in which species are embedded in a low to moderate-dimensional latent space. The challenge to overcome is that such embeddings are typically “learned”, which would result in a nonlinear parameter estimation.
5 Acknowledgements
The work was supported by the New Frontiers in Research Fund through an NFRF Exploration Grant, and by the Natural Sciences and Engineering Research Council of Canada through an NSERC Discovery grant and an NSERC-NSF Collaboration on quantum science and artificial intelligence grant. CO is a partner in Symmetric Group LLP which licenses force fields commercially. The authors acknowledge helpful discussions with Gábor Csányi and Cas van der Oord. This research was supported in part through computational resources and services provided by Advanced Research Computing at the University of British Columbia.
6 Appendix
6.1 Radial Basis Functions
@@ 6.2 The BPDual Algorithm
[…]
@@
The BPDual algorithm is an active-set method. The active set method focuses on identifying constraints that are active (i.e., binding) at each iteration. For an estimated solution y, a constraint is active if $\textbf{a}_{j}^{\top}\textbf{y}$ lies at one of its bounds (i.e. $\textbf{a}_{j}^{\top}\textbf{y}=\mathbf{1}$ or $\mathbf{1}$). The active set $S$ is a set of indices representing the active constraints at each iteration, which forms a submatrix of A that has full column rank. For instance, for some permuation $\mathrm{P}$
@@
The estimate y is stationary if it satisfies the constraints $\mathbf{1}\leq\textbf{A}^{\top}\textbf{y}\leq\mathbf{1}$ and the objective gradient $\mathbf{g}(\mathbf{y})$ is a linear combination of active constraint gradients. Specifically, there exists $x_{S}$ such that:
@@
These conditions imply that the current estimates $(\textbf{x},\textbf{y})$ satisfy feasibility for the primal problem, with x residing within the active constraints.
@@
An overview of the BPDual algorithm is provided in Figure 9. It’s not worthy that for efficient computations, only the upper triangular part of a QR factorization of $S$ is updated at each iteration using the QRupdate.jl [31] package.
Figure 9: A schematic of the BPDual algorithm.
@@
When $\alpha=1$, the solution $\textbf{y}+\Delta\textbf{y}$ minimizes $\phi(\textbf{y})$ on the current active set $S$. If improvement is possible by removing a constraint from $S$, the algorithm considers directions $\Delta\textbf{y}$ that move y away from an active constraint, maintaining feasibility. Consider an inequality constraint $q=S_{p}$ that is currently active at its lower bound, i.e., $a_{q}^{T}\textbf{y}=-\mathbf{1}$. To move away from this bound, we select a direction $\Delta\textbf{y}$ satisfying $S^{T}\Delta\textbf{y}=e_{p}$, ensuring $\Delta\textbf{y}$ is a feasible direction that moves away from the active lower bound. Observe that:
@@
Similarly, if the constraint $q$ is active at its upper bound and $(\mathbf{x}_{\mathcal{S}})_{p}<0$, we can improve the objective by moving in a direction $\Delta\textbf{y}$ for which $S^{T}\Delta\textbf{y}=-e_{p}$. In this case, we also remove constraint $q$ from the active set. Hence, the Lagrange multipliers $\mathbf{x}_{\mathcal{S}}$ indicate which indices in the active set $S$ should be removed to enable a reduction in the objective function. If none of the elements in $\mathbf{x}_{\mathcal{S}}$ have the appropriate sign, then the current point $(\textbf{x},\textbf{y})$ is optimal.
6.3 Lasso Path
@@
As $\lambda$ decreases to $\bar{\lambda}$, it is also necessary to assess the impact on $\mathbf{y}$, particularly with respect to the constraints $\mathbf{1}\leq A^{T}\mathbf{y}\leq\mathbf{1}$. Subtracting $\bar{\mathbf{y}}$ from both sides and using the definition of $\Delta\mathbf{x}$, we obtain:
@@
6.4 Orthogonal Matching Pursuit
Orthogonal Matching Pursuit (OMP) is a greedy algorithm for finding a sparse solution to the linear system $Ax=b$. OMP constructs the support of $x$ iteratively, adding one nonzero component at each step. Initially, the solution vector is set to $x_{0}=0$, the active set is $\mathcal{S}_{0}=\emptyset$, and the residual is $r_{0}=b$. At each iteration $k$, OMP identifies the column $a_{i}$ of $A$ that is most correlated with the current residual $r_{k-1}$. Specifically, it selects the index $i$ that maximizes the absolute value of the following inner product, $i=\underset{i}{\arg\max}|\langle a_{i},r_{k-1}\rangle|$. The active set is then updated as $\mathcal{S}_{k}=\mathcal{S}_{k-1}\cup\{i_{k}\}$. The algorithm computes the new solution $x_{k}$ by minimizing the residual norm $\|b-Ax\|_{2}$ over vectors supported on $S_{k}$. Finally, the residual is updated as $r_{k}=b-Ax_{k}$. The algorithm enforces the orthogonality of $r_{k}$ with respect to the columns of $A$ in the current support, ensuring that once a column is selected, it will not be chosen again. This process repeats until a stopping criterion, such as a sufficiently small residual norm, is met.
6.5 BRR Posterior Derivation
@@
$\boldsymbol{\Sigma}=(\alpha\mathbf{I}+\lambda A^{\top}A)^{-1},$
@@
$\ln p(b\mid A,\alpha,\lambda)=$ $-\frac{m}{2}\ln(2\pi)+\frac{1}{2}\ln|\boldsymbol{\Sigma}|$
@@
This marginal log-likelihood is maximized with respect to $\alpha$ (or $\boldsymbol{\alpha}$ in the ARD case) and $\lambda$, typically using optimization methods such as BFGS or L-BFGS.
6.6 Truncated Singular Value Decomposition (TSVD) Post-Processing